# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

### v3

@@ Abstract
In-Context Operator Networks (ICONs) have demonstrated the ability to learn operators across diverse partial differential equations using few-shot, in-context learning. However, existing ICONs process each spatial point as an individual token, severely limiting computational efficiency when handling dense data in higher spatial dimensions. We propose Vision In-Context Operator Networks (VICON), which integrates vision transformer architectures to efficiently process 2D data through patch-wise operations while preserving ICON’s adaptability to multiphysics systems and varying timesteps. Evaluated across three fluid dynamics benchmarks, VICON significantly outperforms state-of-the-art baselines: DPOT and MPP, reducing the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times. VICON naturally supports flexible rollout strategies with varying timestep strides, enabling immediate deployment in imperfect measurement systems where sampling frequencies may differ or frames might be dropped—common challenges in real-world settings—without requiring retraining or interpolation. In these realistic scenarios, VICON exhibits remarkable robustness, experiencing only 24.41% relative performance degradation compared to 71.37%-74.49% degradation in baseline methods, demonstrating its versatility for depolying in realistic applications. Our scripts for processing datasets and code are publicly available at https://github.com/Eydcao/VICON.
@@ 1 Introduction
This limitation led to the development of “operator learning,” where neural networks learn mappings from input functions to output functions, such as from initial/boundary conditions to PDE solutions. This capability was first demonstrated using shallow neural networks [8, 9], followed by specialized architectures including the widely-adopted Deep Operator Network (DeepONet) [10], Fourier Neural Operator (FNO) [11, 12], and recent transformer-based operator learners [13, 14, 15]. While these methods excel at learning individual parametric PDEs, they cannot generalize across different PDE types, necessitating costly retraining for new equations or even different time step sizes.
@@
To further lower barriers for applying multi-physics models in real-world applications, the In-Context Operator Network (ICON) [25] offers a fundamentally different approach: ICON implicitly encodes system dynamics through a few input-output function pairs, then extracts dynamics from these pairs in an in-context fashion. This approach enables few-shot generalization without retraining or explicit PDE knowledge, while the minimal required input-output pairs can be readily collected during deployment. A single ICON model has demonstrated success in handling both forward and inverse problems across various ODEs, PDEs, and mean field control scenarios.
@@ 2 Related Works
Operator learning [8, 9, 11, 10] addresses the challenge of approximating operators $G:U\to V$, where $U$ and $V$ are function spaces representing physical systems and differential equations, e.g., $G$ maps initial/boundary conditions or system parameters to corresponding solutions. Among the various operator learning approaches, DeepONets [10, 28] employ branch and trunk networks to independently process inputs and query points, while FNOs [11] leverage fast Fourier transforms to efficiently compute kernel integrations for PDE solutions in regular domains. These methods have been extended to incorporate equation information [29, 30], multiscale features [31, 32, 33], adaptation to heterogeneous and irregular meshes [34, 35, 36, 37].
However, these approaches typically learn a single operator, requiring retraining when encountering different PDE types or time-step sizes. Our work aims to develop a unified model for multiple PDEs with few-shot generalization capabilities based on limited observation frames.
@@
Foundation models in natural language processing [38, 39] and computer vision [40] have demonstrated remarkable versatility across diverse tasks. Drawing inspiration from this paradigm, recent research has explored developing unified models for multiple PDEs in domains such as PDE discovery [41] and computational fluid dynamics [42]. Several approaches follow pre-training and fine-tuning strategies [16, 17], though these require additional resources for data acquisition and fine-tuning before deployment. In parallel, researchers have developed zero- or few-shot multi-physics models by incorporating additional information; for instance, the PROSE approach [19, 20, 21, 22, 43] directly encodes symbolic information of PDEs into the model. Other methods achieving zero- or few-shot generalization include using physics-informed tokens [24], representing PDE structures as graphs [23], or conditioning transformers on PDE descriptions [44]; however, all these techniques require prior knowledge of the PDEs as additional inputs. We select two state-of-the-art models, MPP [26] and DPOT [27], as our baselines.
@@ 3 Preliminaries
Denote an ICON model as $\mathcal{T}_{\theta}$, where $\mathcal{T}$ is a transformer with trainable parameters $\theta$. The model takes input as a sequence comprising $I$ pairs of conditions (${\bm{c}}$) and quantities of interest (${\bm{q}}$) as $\{\langle{\bm{c}}_{i},{\bm{q}}_{i}\rangle\}_{i=1}^{I}$, where each ${\bm{c}},{\bm{q}}$ contains multiple tokens representing a single function. The model outputs tokens representing future functions at the positions of ${\bm{c}}$, i.e. “next function prediction” similar to “next token predictions” in LLM. More precisely, for $J\in\{1,\dots,I-1\}$, the model predicts $\tilde{{\bm{q}}}_{J+1}$ given the leading $J$ pairs and the condition ${\bm{c}}_{J+1}$:
@@
After training, ICON can process a new set of $\langle{\bm{c}},{\bm{q}}\rangle$ pairs in the forward pass to in-contextually learn operators and employ them to predict future functions using new conditions:
@@
where $J\in\{I_{\text{min}},\dots,I-1\}$ is the number of in-context examples, $I_{\text{min}}$ is the number of context exempted from loss calculation, and $k>I$ is the index of the condition for future prediction.
Importantly, all pairs in the same sequence must be formed with the same operator mapping (i.e., the same PDE and consistent timestep size), while operators can vary across different sequences.
@@ 4.2 Vision In-Context Operator Networks
Inspired by the Vision Transformer (ViT) [47], we address these limitations by dividing the physical fields into patches, where each patch is flattened and projected as a token. This approach, which we call Vision In-Context Operator Networks (VICON), has its forward process illustrated in Figure 1(b) and detailed in the following.
@@
where $k=1,\ldots,N_{c}\text{ and }l=1,\ldots,N_{q}\text{ are the index of patches.}$.
We inject two types of learnable positional encoding before feeding the embeddings into the transformer: (1) patch positional encodings to indicate relative patch positions inside the whole domain, denoted as $\mathbf{E}_{p}\in\mathbb{R}^{N_{p}\times d}$, where $N_{p}=\max\{N_{c},N_{q}\}$; (2) function positional encodings to indicate whether a patch belongs to a input function (using $\mathbf{E}_{c}\in\mathbb{R}^{I\times d}$) or output function (using $\mathbf{E}_{q}\in\mathbb{R}^{I\times d}$), as well as their indices in the sequence:
@@
To support autoregressive prediction similar to Equation 2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models [38] and large vision models [48]. The mask is defined as follows:
@@
After obtaining the output tokens from Equation 2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@ 4.3 Prompt Normalization
To address the varying scales across different channels and prompts (i.e. sequences of pairs $\{\langle{\bm{c}}_{i},{\bm{q}}_{i}\rangle\}_{i=1}^{I}$), we normalize the data before feeding them into the model. A crucial requirement is to consistently normalize functions within the same sequence, to ensure that the same operator is learned. Denoting the normalization operators for ${\bm{c}}$ and ${\bm{q}}$ as $\mathcal{N}_{c}$ and $\mathcal{N}_{q}$ respectively, and the operator in the original space as $\mathcal{L}$, the operator in the normalized space $\mathcal{L}^{\prime}$ follows:
@@ 4.4 Datasets and Data Augmentation
We evaluate on three fluid dynamics datasets representing different physical regimes: 1) PDEArena-Incomp [1] (incompressible Navier-Stokes equations), containing 2,496/608/608 trajectories (train/valid/test) with 56 timesteps each; 2) PDEBench-Comp-HighVis [49] (compressible Navier-Stokes with high viscosity), containing 40,000 trajectories of 21 timesteps each; and 3) PDEBench-Comp-LowVis [49] (compressible Navier-Stokes with numerically zero viscosity), containing 4,000 trajectories of 21 timesteps each. For PDEBench-Comp-HighVis and PDEBench-Comp-LowVis, we randomly split trajectories in 80%/10%/10% proportions for training/validation/testing, respectively.
The temporally homogeneous Markovian property (Section 4.1) enables natural data augmentation for training and flexible rollout strategy for inference (Section 4.5): by striding with larger timesteps to reduce the number of autoregressive steps in long-term prediction tasks: $\Delta t=s\Delta\tau_{i_{p}}\;|\;s=1,2,\dots,s_{\text{max}}$, where $\tau_{i_{p}}$ is the timestep size for recording in trajectory $i_{p}$. During training, for each trajectory in the training set, we first sample a stride size $s\sim\mathcal{U}\{1,\ldots,s_{\text{max}}\}$, then form $\langle\mathbf{u}_{t},\mathbf{u}_{t+s\Delta\tau}\rangle$ pairs for the corresponding operator. We illustrate this augmentation process in Figure 1(a), where different strides are randomly sampled during dataloading.
Additional dataset details appear in Appendix A.
@@ 4.5 Inference with Flexible Strategies
As mentioned in Section 4.4, VICON can make predictions with varying timestep strides. Given $I_{0}$ initial frames $\{\mathbf{u}_{i}\}_{i=1}^{I_{0}}$, we can form in-context example pairs $\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+s}}\rangle\}_{i=1}^{I_{0}-s}$. For $j\geq 1$, we set the question condition (${\bm{c}}_{k}$ in Equation 3) as ${\bm{u}}_{j}$ to predict ${\bm{u}}_{j+s}$, enabling $s$-step prediction.
@@
This strategy is also applicable to the imperfect measuring cases where partial input frames are missing, eg, due to sensor device error. In this case, we can still form pairs from the remaining frames, with minor adjustment to the strategy generation algorithms, as shown in Figure 1(c).
More details (including algorithms and examples) on rollout strategies are provided in Section C.1 (for perfect measurements) and in Section C.2 (for imperfect measurements).
@@ 4.6 Evaluation Metric
We evaluate the rollout accuracy of VICON using two different strategies described in Section 4.5. The evaluation uses relative and absolute $L^{2}$ errors between the predicted and ground truth frames, starting from the frame $I_{0}+1$. For relative scaling coefficients, we use channel-wise scaling standard deviation $\mathbf{\sigma}$ of ground truth frames, which vary between different prompts.
@@ 5 Experimental Results
For evaluation, we use both absolute and relative $L^{2}$ RMSE (Section 4.6) on rollout predictions. Since VICON offers the flexibility of predictions with different time step strides, we evaluate our single trained model using both single-step and flexible-step rollout strategies (Section 4.5) during inference.
For conciseness, we present summarized plots and tables the main text, and defer complete results and visualizations to Appendix D. Ablation studies examining key design choices of VICON—including patch resolution, positional encoding, and context length—are presented in Section D.1.
@@
Last step [1e-2] PDEArena-Incomp 76.77 68.03 65.27 93.52
@@
All average [1e-2] PDEArena-Incomp 56.27 48.50 41.20 55.95
@@ 5.1 Superior Performance on Long-term Rollout Predictions
As demonstrated in Figure 2(a), Table 1, and Figure 3, VICON consistently outperforms baseline methods on long-horizon predictions across all benchmarks—with the exception of DPOT on the PDEArena-Incomp dataset, where our performance is comparable. Overall, VICON achieves an average reduction in relative $L^{2}$ RMSE at the final timestep of 37.9% compared to DPOT and 44.7% compared to MPP.
Notably, DPOT exhibits exceptionally poor performance on the PDEBench-Comp-HighVis dataset, with an 821.9% error compared to our 30.06%. Our visualization of failure cases reveals that DPOT struggles with trajectories with small pressure values compared to the dataset average. This may stems from DPOT’s lack of prompt normalization, as compressible flow’s pressure channel exhibit large magnitude variations.
@@
More detailed results are provided in Table 8 and Table 9 in Appendix D.

### v4

@@ Abstract
In-Context Operator Networks (ICONs) have demonstrated the ability to learn operators across diverse partial differential equations using few-shot, in-context learning. However, existing ICONs process each spatial point as an individual token, severely limiting computational efficiency when handling dense data in higher spatial dimensions. We propose Vision In-Context Operator Networks (VICON), which integrate vision transformer architectures to efficiently process 2D data through patch-wise operations while preserving ICON’s adaptability to multi-physics systems and varying timesteps. Evaluated across three fluid dynamics benchmarks, VICON significantly outperforms state-of-the-art baselines DPOT and MPP, reducing the average last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times. VICON naturally supports flexible rollout strategies with varying timestep strides, enabling immediate deployment in imperfect measurement systems where sampling frequencies may differ or frames might be dropped—common challenges in real-world settings—without requiring retraining or interpolation. In these realistic scenarios, VICON exhibits remarkable robustness, experiencing only 24.41% relative performance degradation compared to 71.37%-74.49% degradation in baseline methods, demonstrating its versatility for deployment in realistic applications. Our scripts for processing datasets and code are publicly available at https://github.com/Eydcao/VICON.
@@ 1 Introduction
This limitation led to the development of “operator learning,” where neural networks learn mappings from input functions to output functions, such as from initial/boundary conditions to PDE solutions. This capability was first demonstrated using shallow neural networks [8, 9], followed by specialized architectures including the widely adopted Deep Operator Network (DeepONet) [10], Fourier Neural Operator (FNO) [11, 12], and recent transformer-based operator learners [13, 14, 15]. While these methods excel at learning individual parametric PDEs, they cannot generalize across different PDE types, necessitating costly retraining for new equations or even different timestep sizes.
@@
To further lower barriers for applying multi-physics models in real-world applications, the In-Context Operator Network (ICON) [25] offers a fundamentally different approach: ICON implicitly encodes system dynamics through a few input-output function pairs, then extracts dynamics from these pairs in an in-context fashion. This approach enables few-shot generalization without retraining or explicit PDE knowledge, while the minimal required input-output pairs can be readily collected during deployment. A single ICON model has demonstrated success in handling both forward and inverse problems across various ODEs, PDEs, and mean-field control scenarios.
@@ 2 Related Work
Operator learning [8, 9, 11, 10] addresses the challenge of approximating operators $G:U\to V$, where $U$ and $V$ are function spaces representing physical systems and differential equations, e.g., $G$ maps initial/boundary conditions or system parameters to corresponding solutions. Among the various operator learning approaches, DeepONets [10, 28] employ branch and trunk networks to independently process inputs and query points, while FNOs [11] leverage fast Fourier transforms to efficiently compute kernel integrations for PDE solutions in regular domains. These methods have been extended to incorporate equation information [29, 30], multiscale features [31, 32, 33], and adaptation to heterogeneous and irregular meshes [34, 35, 36, 37].
However, these approaches typically learn a single operator, requiring retraining when encountering different PDE types or timestep sizes. Our work aims to develop a unified model for multiple PDEs with few-shot generalization capabilities based on limited observation frames.
@@
Foundation models in natural language processing [38, 39] and computer vision [40] have demonstrated remarkable versatility across diverse tasks. Drawing inspiration from this paradigm, recent research has explored developing unified models for multiple PDEs in domains such as PDE discovery [41] and computational fluid dynamics [42]. Several approaches follow pre-training and fine-tuning strategies [16, 17], though these require additional resources for data acquisition and fine-tuning before deployment. In parallel, researchers have developed zero- or few-shot multi-physics models by incorporating additional information; for instance, the PROSE approach [19, 20, 21, 22, 43] directly encodes symbolic information of PDEs into the model. Other methods achieving zero- or few-shot generalization include using physics-informed tokens [24], representing PDE structures as graphs [23], or conditioning transformers on PDE descriptions [44]. However, all these techniques require prior knowledge of the PDEs as additional inputs. We select two state-of-the-art models, MPP [26] and DPOT [27], as our baselines.
@@ 3 Preliminaries
Denote an ICON model as $\mathcal{T}_{\theta}$, where $\mathcal{T}$ is a transformer with trainable parameters $\theta$. The model takes input as a sequence comprising $I$ pairs of conditions ($\bm{c}$}) and quantities of interest ($\bm{q}$}) as $\{\langle{\bm{c}}_{i},{\bm{q}}_{i}\rangle\}_{i=1}^{I}$, where each ${\bm{c}},{\bm{q}}$ contains multiple tokens representing a single function. The model outputs tokens representing future functions at the positions of $\bm{c}$}, i.e., “next function prediction” similar to “next token predictions” in LLMs. More precisely, for $J\in\{1,\dots,I-1\}$, the model predicts $\tilde{{\bm{q}}}_{J+1}$ given the leading $J$ pairs and the condition ${\bm{c}}_{J+1}$:
@@
After training, ICON can process a new set of $\langle{\bm{c}},{\bm{q}}\rangle$ pairs in the forward pass to learn operators in-context and employ them to predict future functions using new conditions:
@@
where $J\in\{I_{\text{min}},\dots,I-1\}$ is the number of in-context examples, $I_{\text{min}}$ is the number of context examples exempted from loss calculation, and $k>I$ is the index of the condition for future prediction.
Importantly, all pairs in the same sequence must be formed with the same operator mapping (i.e., the same PDE and consistent timestep size within one sequence), while operators can vary across different sequences/rows in the training batch to enable the model’s generalization ability to different combinations of PDEs and timestep sizes.
@@ 4.2 Vision In-Context Operator Networks
Inspired by the Vision Transformer (ViT) [47], we address these limitations by dividing the physical fields into patches, where each patch is flattened and projected as a token. This approach, which we call Vision In-Context Operator Networks (VICON), addresses the computational limitation of treating each spatial point as an individual token, which leads to quadratic computational complexity with respect to the number of points. This patch-wise approach has its forward process illustrated in fig.˜1(b) and detailed in the following.
@@
where $k=1,\ldots,N_{c}\text{ and }l=1,\ldots,N_{q}\text{ are the index of patches.}$
We inject two types of learnable positional encoding before feeding the embeddings into the transformer: (1) patch positional encodings to indicate relative patch positions inside the whole domain, denoted as $\mathbf{E}_{p}\in\mathbb{R}^{N_{p}\times d}$, where $N_{p}=\max\{N_{c},N_{q}\}$ and the encoding varies per patch (broadcast across points within each patch); (2) function positional encodings to indicate whether a patch belongs to a input function (using $\mathbf{E}_{c}\in\mathbb{R}^{I\times d}$) or output function (using $\mathbf{E}_{q}\in\mathbb{R}^{I\times d}$), as well as their indices in the sequence, where the encoding varies per condition/QoI in each pair (broadcast across patches and points of that condition/QoI):
@@
To support autoregressive prediction similar to eq.˜2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models [38] and large vision models [48]. The mask is defined as follows:
@@
After obtaining the output tokens from eq.˜2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@ 4.3 Prompt Normalization
To address the varying scales across different channels and prompts (i.e., sequences of pairs $\{\langle{\bm{c}}_{i},{\bm{q}}_{i}\rangle\}_{i=1}^{I}$), we normalize the data before feeding them into the model. A crucial requirement is to consistently normalize functions within the same sequence, to ensure that the same operator is learned. Denoting the normalization operators for $\bm{c}$} and $\bm{q}$} as $\mathcal{N}_{c}$ and $\mathcal{N}_{q}$ respectively, and the operator in the original space as $\mathcal{L}$, the operator in the normalized space $\mathcal{L}^{\prime}$ follows:
@@ 4.4 Datasets and Data Augmentation
We evaluate on three fluid dynamics datasets representing different physical regimes: 1) PDEArena-Incomp [1] (incompressible Navier-Stokes equations), containing 2,496/608/608 trajectories (train/valid/test) with 56 timesteps each; 2) PDEBench-Comp-HighVis [49] (compressible Navier-Stokes with high viscosity), containing 40,000 trajectories of 21 timesteps each; and 3) PDEBench-Comp-LowVis [49] (compressible Navier-Stokes with numerically zero viscosity), containing 4,000 trajectories of 21 timesteps each. For PDEArena-Incomp, we follow the original data split in [1]. For PDEBench-Comp-HighVis and PDEBench-Comp-LowVis, we randomly split trajectories in 80%/10%/10% proportions for training/validation/testing, respectively. For all datasets during rollout, we set $I_{0}=10$, i.e., we predict trajectories up to the last timestep given the initial 10 frames.
The temporally homogeneous Markovian property (section˜4.1) enables natural data augmentation for training and flexible rollout strategy for inference (section˜4.5): by striding with larger timesteps to reduce the number of autoregressive steps in long-term prediction tasks: $\Delta t=s\Delta\tau_{i_{p}}\;|\;s=1,2,\dots,s_{\text{max}}$, where $\tau_{i_{p}}$ is the timestep size for recording in trajectory $i_{p}$. During training, for each trajectory in the training set, we first sample a stride size $s\sim\mathcal{U}\{1,\ldots,s_{\text{max}}\}$, then form $\langle\mathbf{u}_{t},\mathbf{u}_{t+s\Delta\tau}\rangle$ pairs for the corresponding operator. We illustrate this augmentation process in fig.˜1(a), where different strides are randomly sampled during dataloading.
Additional dataset details appear in appendix˜A.
@@ 4.5 Inference with Flexible Strategies
As mentioned in section˜4.4, VICON can make predictions with varying timestep strides. Given $I_{0}$ initial frames $\{\mathbf{u}_{i}\}_{i=1}^{I_{0}}$, we can form in-context example pairs $\{\langle\mathbf{u}_{i},\mathbf{u}_{{i+s}}\rangle\}_{i=1}^{I_{0}-s}$. For $j\geq 1$, we set the question condition (${\bm{c}}_{k}$ in eq.˜3) as ${\bm{u}}_{j}$ to predict ${\bm{u}}_{j+s}$, enabling $s$-step prediction.
@@
This strategy is also applicable to imperfect measurement cases where partial input frames are missing, e.g., due to sensor device error. In this case, we can still form pairs from the remaining frames, with minor adjustments to the strategy generation algorithms, as shown in fig.˜1(c).
More details (including algorithms and examples) on rollout strategies are provided in section˜C.1 (for perfect measurements) and in section˜C.2 (for imperfect measurements).
@@ 4.6 Evaluation Metric
We evaluate the rollout accuracy of VICON using two different strategies described in section˜4.5. The evaluation uses relative and absolute $L^{2}$ errors between the predicted and ground truth frames, starting from the frame $I_{0}+1$. For relative scaling coefficients, we use channel-wise scaling standard deviation $\mathbf{\sigma}$ of ground truth frames, which vary between different prompts.
@@ 5 Experimental Results
For evaluation, we use both absolute and relative $L^{2}$ RMSE (section˜4.6) on rollout predictions. Since VICON offers the flexibility of predictions with different timestep strides, we evaluate our single trained model using both single-step and flexible-step rollout strategies (section˜4.5) during inference.
For conciseness, we present summarized plots and tables in the main text, and defer complete results and visualizations to appendix˜D. Ablation studies examining key design choices of VICON—including patch resolution, positional encoding, and context length—are presented in section˜D.1. The checkerboard artifacts (e.g., in fig.˜9) of the PDEBench-Comp-HighVis dataset are discussed in section˜D.2.
@@
Last step
[1e-2]
PDEArena-Incomp 76.77 68.03 65.27 93.52
@@
All average
[1e-2]
PDEArena-Incomp 56.27 48.50 41.20 55.95
@@ 5.1 Superior Performance on Long-term Rollout Predictions
As demonstrated in fig.˜2(a), table˜1, and fig.˜3, VICON consistently outperforms baseline methods on long-horizon predictions across all benchmarks—with the exception of DPOT on the PDEArena-Incomp dataset, where our performance is comparable. Overall, VICON achieves an average reduction in relative $L^{2}$ RMSE at the final timestep of 37.9% compared to DPOT and 44.7% compared to MPP.
Notably, DPOT exhibits exceptionally poor performance on PDEBench-Comp-HighVis, with an 821.9% error compared to our 30.06%. Our visualization of failure cases reveals that DPOT struggles with trajectories with small pressure values compared to the dataset average. This may stem from DPOT’s lack of prompt normalization, as compressible flow’s pressure channel exhibits large magnitude variations.
@@
More detailed results are provided in table˜11 and table˜12 in appendix˜D.